find an equation of the line through the given point , write the equation in standard form through (-4,1) and(1,-3)
Answer by Maths68(1474)

Points
(x1, y1) = (-4,1)
(x2, y2) = (1, -3)
Use Point-slope form
m=(y2-y1)/(x2-x1)
Plug in the values in above equation
m=(-3-1)/(1-(-4))
m=(-4)/(1+4)
m=-4/5
Use equation of slope-intercept form
y=mx+b
Put the values of m and any of two points (-4,1) or (1,-3)in above equation
first (-4, 1)
1=(-4/5)(-4)+b
1=16/5 + b
1 - 16/5 = b
5-16 / 5 = b
-11/5=b
b=-11/5

Now (1,-3)
y=mx+b
-3=(-4/5)(1)+b
-3=-4/5+b
-3 + 4/5 = b
(-15+4)/5 = b
-11/5=b
Put the value of b in the equation of the line (standard form)
y=mx+b
y=(-4/5)x+(-11/5)
y=(-4/5)x-(11/5)